Super unleaded gasoline costs 9 cents more per gallon than regular unleaded gasoline. A service station buys 3,000 gallons of super unleaded and 2,000 gallons of regular unleaded for $8,919. What is the cost per gallon of each kind of gasoline?
Answer by mananth(16949)

unleaded = x /gallon
super ---- (x+9)

2000x+3000(x+9)=8919
2000x+3000x+27000=891900
5000x=891900-27000
5000x=864900
x=172.98 cents/gallon
x= $1.73/gallon unleaded
super = $1.82 / gallon